# Faster-than-Nyquist Signaling in the Finite Time-Bandwidth Product Regime

### v3

@@ Appendix C Approximation Of FTN MC and RCU Bounds
Consider a linear combination $Y=\sum_{n=0}^{N-1}X_{n}$, where $X_{n}$ with PDF $f_{X_{n}}(x)$ are independent but not necessarily identically distributed. Let $K_{Y}(t)=\ln\mathbb{E}\{e^{tY}\}$ be the cumulative generating function (CGF) of $Y$ and $K^{\prime}_{Y}(t)$ and $K^{\prime\prime}_{Y}(t)$ be 11st and 22nd derivatives of $K_{Y}(t)$, respectively. Then, due to Corollary 1 of [maya2016closed], the logarithm of the CDF of $Y$ may be approximated as
@@ Appendix E PSWF: Definitions, properties, and benchmarks
Let $\phi_{\Omega,n}(t)\triangleq\frac{\mathcal{D}\psi_{\Omega,n}(t)}{\sqrt{\mu_{\Omega,n}}}$ denote normalized and truncated PSWFs with the time support $|t|\leq\frac{T_{x}}{2}$ and a unit energy. These are orthonormal and complete in the space of time-limited functions in $|t|\leq T_{x}/2$. Furthermore, its energy concentration within the band $|f|\leq W$ is also $\mu_{\Omega,n}$, i.e., $\int_{-W}^{W}{|\hat{\phi}_{\Omega,n}(f)|^{2}df}=\mu_{\Omega,n}$, where $\hat{\phi}_{\Omega,n}(f)$ is the Fourier transform of $\phi_{\Omega,n}(t)$.

### v4

@@ Appendix C Approximation Of FTN MC and RCU Bounds
Consider a linear combination $Y=\sum_{n=0}^{N-1}X_{n}$, where $X_{n}$ with PDF $f_{X_{n}}(x)$ are independent but not necessarily identically distributed. Let $K_{Y}(t)=\ln\mathbb{E}\{e^{tY}\}$ be the cumulative generating function (CGF) of $Y$ and $K^{\prime}_{Y}(t)$ and $K^{\prime\prime}_{Y}(t)$ be the $1^{\text{st}}$ and $2^{\text{nd}}$ derivatives of $K_{Y}(t)$, respectively. Then, due to Corollary 1 of [maya2016closed], the logarithm of the CDF of $Y$ may be approximated as
@@ Appendix E PSWF: Definitions, properties, and benchmarks
Let $\phi_{\Omega,n}(t)\triangleq\frac{\mathcal{D}\psi_{\Omega,n}(t)}{\sqrt{\mu_{\Omega,n}}}$ denote normalized and time-truncated PSWFs with the time support $|t|\leq\frac{T_{x}}{2}$ and a unit energy. These are orthonormal and complete in the space of time-limited functions in $|t|\leq T_{x}/2$. Furthermore, its energy concentration within the band $|f|\leq W$ is also $\mu_{\Omega,n}$, i.e., $\int_{-W}^{W}{|\hat{\phi}_{\Omega,n}(f)|^{2}df}=\mu_{\Omega,n}$, where $\hat{\phi}_{\Omega,n}(f)$ is the Fourier transform of $\phi_{\Omega,n}(t)$.